y varies directly with x, y=65 when x=5. find y when x=12.
Answer by Theo(13342)

the formula for for y varies directly with x is:

y = k*x

when y = 65, x = 5, so your equation becomes:

65 = 5*k

solve for k and you get k = 13 because 13 * 5 = 65.

question was find y when x = 12.

equation of y = k*x becomes y = 13 * 12 = 156.

you have y = 156 when x = 13.

we solved this using the formula y = k*x.

we could also have solved this by using the ratio equation of:

5/65 = 12/y

cross multiply to get 5*y = 12*65

divide both sides of this equation by 5 to get:

y = 12*65/5 which becomes y = 12*13 which becomes y = 156

This website has the formulas for variation that you would need.

Direct and Inverse Variation Formulas plus Others